STRUCTURE AND SPECTROSCOPIC CONSTANTS OF FeCN: AN AB INITIO MOLECULAR ORBITAL STUDY

\underline{TSUNEO HIRANO}, YUKARI MITSUI , Department of Chemistry, Faculty of Science, Ochanomizu University, 2-2-1 Otsuka, Bunkyo-ku, Tokyo 112-8610, Japan; UMPEI NAGASHIMA, National Institute of Materials and Chemical Research, 1-1 Higashi, Tsukuba, Ibaraki 305-8565, Japan.

FeCN is one of the possible candidates for an Fe-bearing interstellar molecule. Since no experimental data is available, prediction of its structure and spectroscopic constants by ab initio molecular orbital (MO) calculations has been requested. The ground state FeCl is known to be ^{6}Δ, and Cl^{−} and CN^{−} are known to show a similar chemical behavior. Hence, we assumed ^{6}Δ state for the ground state FeCN. Wachters’ all-electron basis set augmented with Bauschlicher’s f-GTO, [8s,6p,4d,1f], for Fe and Dunning’s cc-pVTZ basis sets for C and N have been employed. Active space for the MR-SDCI calculations has been selected to include Fe 3d and 4s as well as CN π and π^{∗} orbitals obtained by the full valence MC-SCF MO calculations. Three dimentional potential energy surface has been calculated at the MR-SDCI level, and analyzed by the Mills’ 2nd order perturbation theory. The ^{6}Δ FeCN is predicted to be a linear molecule with Fe-C and CN bond lengths of 2.066 and 1.169 Å, respectively, yielding B_{e}, D_{J}, and B_{0} to be 3667.9, 0.0012, and 3676.1 MHz. Anharmonic vibrational frequencies ν_{1}, ν_{2}, and ν_{3} are predicted to be 2158, 185, and 450 cm^{−1}, respectively.